OpenStudy (anonymous):

A cinema ticket for one adult and one child costs £13.50.The cost of 2 adults and 5 children is £24 find the cost of tickets for one adult and one child

OpenStudy (anonymous):

x+y=13.50 2x+5y=24 now multiply equation one by 2 2x+2y=27 2x+5y=24 subtract equation 2 from 1 yor gonna get 3y=3 y=1 now put y=1 in equation 1 yor gona get x+1=13.50 x=12.50 answers :)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

hope that helps ya chloeeeee